Heliox, a breathing gas composed of a mixture of helium (He) and oxygen (O₂), is used medically for treatment in acute asthma. In medicine heliox may refer to a mixture of 21% O2 (the same as air) and 79% He. Calculate the average molar mass and density of heliox at pressure Patm = 100 kPa and room temperature t = 22°C. The molar mass of oxygen is po 16 and helium He = 4 g/mole.


1
Expert's answer
2021-05-20T10:11:59-0400

Solution.

"m(O_2)=0.21m;"

"m(He)=0.79m;"

"P_{atm}=100\sdot10^3Pa;"

"T=295K;"

"M(O_2)=32\sdot10^{-3}kg/mol;"

"M(He)=4\sdot10^{-3}kg/mol;"

"\nu=\nu(O_2)+\nu(He);"

"\nu=\dfrac{0.21m}{32\sdot10^{-3}}+\dfrac{0.79m}{4\sdot10^{-3}}=\dfrac{6,53m}{32\sdot10^{-3}}=2.04\sdot10^2(mol);"


"\nu=\dfrac{m}{M}\implies M=\dfrac{m}{\nu};"

"M=\dfrac{m}{2.04\sdot10^2}=5\sdot10^{-3}(kg/mol);"

"PV=\dfrac{m}{M}RT; P=\dfrac{\rho}{M}RT\implies \rho=\dfrac{PM}{RT};"


"\rho=\dfrac{100\sdot10^3\sdot5\sdot10^{-3}}{8.31\sdot295}=2.04(kg/m^3);"

Answer:"M\n\n\u200b\t\n =5\u22c510^\n{\u22123}\n (kg\/mol);"

"\rho=2.04(kg/m^3)."